# Probing LLMs for Multilingual Discourse Generalization Through a Unified Label Set

### v1

@@ Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, especially to Philipp Mondorf, Siyao Peng, and Shijia Zhou. We would also like to express our gratitude to Amir Zeldes for the constructive and valuable feedback. Lastly, we recognize the support for Yang Janet Liu and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
@@ Appendix A Unified Label Set: Definitions and Examples
informs about the source of information and is considered a discourse relation by RST and SDRT. PDTB’s attribution annotation is considered a separate discourse annotation type and is not included in DISRPT.
@@ Appendix C Models and Probe Training
For each model, we compute one pass over the dataset computing all the attention representations which we cache for the probing experiments. For the smaller models, a cluster equipped with eight Nvidia A100 GPUs was used for around $80$ hours. For the large models of size $70$B and $72$B, we used a cluster equpped with four Nvidia H200 GPUs for about $30$ hours.
@@
As specified in the code repository, we release our code under MIT license.
@@
Table 2: Overview of the included LLMs and their multilingual capabilities for languages in DISRPT.
@@
Table 3: Dataset accuracy scores by LLM probe averaged over five runs. Last two columns refer to the DisCoDisCo reference system trained on all languages (all) and trained a multiple models with different encoders per dataset.
Appendix D Additional Results and Analysis
Figure 6 shows the layer-wise sums of the dot-products of the weight matrices. The magnitude of these scores can be interpreted as a feature importance and confirm that earlier layers play a crucial role in predictions, although the higher layers also show higher scores from layers $30$ to $38$, indicating a progressive refinement of discourse representations. This is in accordance with previous literature, which has shown that the lower layers of LLMs are mostly involved with encoding, while the upper layers are involved in generating and updating the next token prediction Ju et al. (2024).

### v2

@@ Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, especially to Philipp Mondorf, Siyao Peng, and Shijia Zhou. We would also like to express our gratitude to Amir Zeldes for the constructive and valuable feedback. We thank the reviewers and the meta-reviewer for their feedback and suggestions. Lastly, we recognize the support for Yang Janet Liu and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
@@ Appendix A Unified Label Set: Definitions and Examples
informs about the source of information, which is useful and crucial for many real-world applications such as misinformation detection and fact-checking. It is considered a discourse relation by RST, SDRT, and DEP. PDTB’s attribution annotation is considered a separate discourse annotation type and is not included in DISRPT. It is worth pointing out that attribution is not considered a core discourse relation by Bunt and Prasad (2016), but it is one of the most frequent discourse relations in most frameworks.
@@ Appendix C Models and Probe Training
For each model, we compute one pass over the dataset computing all the attention representations which we cache for the probing experiments. For the smaller models, a cluster equipped with eight Nvidia A100 GPUs was used for around $80$ hours. For the large models of size $70$B and $72$B, we used a cluster equipped with four Nvidia H200 GPUs for about $30$ hours.
@@
As specified in the code repository, [^4] https://github.com/mainlp/discourse_probes we release our code under MIT license.
@@
Table 2: Overview of the included LLMs and their multilingual capabilities for languages in DISRPT 2023.
Appendix D Additional Results and Analysis
Figure 7 shows the layer-wise sums of the dot-products of the weight matrices. The magnitude of these scores can be interpreted as a feature importance and confirm that earlier layers play a crucial role in predictions, although the higher layers also show higher scores from layers $30$ to $38$, indicating a progressive refinement of discourse representations.
@@ Appendix D Additional Results and Analysis
Table 3: Dataset accuracy scores by LLM probe averaged over five runs. The last two columns refer to the DisCoDisCo reference system trained on all languages (all) and trained a multiple models with different encoders per dataset.
Llama-3.1-70B
Llama-3.1-8B
Llama-3.2-1B
Llama-3.2-3B
Mistral-Small-24B-Base-2501
Qwen2.5-0.5B
Qwen2.5-1.5B
Qwen2.5-14B
Qwen2.5-32B
Qwen2.5-3B
Qwen2.5-72B
Qwen2.5-7B
aya-23-35B
aya-23-8B
aya-expanse-32b
aya-expanse-8b
bloom-1b1
bloom-3b
bloom-560m
bloom-7b1
bloomz-7b1
emma-500-llama2-7b
phi-4
$\mathbf{C}$ (inter) 53.8 45.5 41.5 43.4 45.5 35.1 35.6 49.1 52.9 40.8 55.8 42.4 53.7 47.6 54.3 47.0 32.4 39.9 32.3 40.3 40.1 44.4 46.9
$\mathbf{D}_{1},\mathbf{D}_{2}$ (intra) 46.0 38.2 35.2 37.3 40.1 33.3 32.9 43.5 44.5 33.4 47.4 38.3 45.5 36.0 39.9 40.3 28.5 30.8 29.0 31.9 31.3 40.4 41.3
$\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}$ (all) 55.9 50.8 46.2 48.6 52.5 41.6 42.2 54.2 55.8 46.0 56.8 50.3 58.2 52.6 56.0 54.0 37.8 42.6 37.0 43.5 43.9 51.4 51.9
Table 4: Overall accuracy scores by LLM probe averaged over five runs training on the entire DISRPT (probe representation ablation). We ablate different types of attention representations used for probing and find that using all described matrices $\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}$ yields the best results.
Llama-3.1-70B
Llama-3.1-8B
Llama-3.2-1B
Llama-3.2-3B
Mistral-Small-24B-Base-2501
Qwen2.5-0.5B
Qwen2.5-1.5B
Qwen2.5-14B
Qwen2.5-32B
Qwen2.5-3B
Qwen2.5-72B
Qwen2.5-7B
aya-23-35B
aya-23-8B
aya-expanse-32b
aya-expanse-8b
bloom-1b1
bloom-3b
bloom-560m
bloom-7b1
bloomz-7b1
emma-500-llama2-7b
phi-4
mean 55.6 48.7 45.2 47.8 48.6 39.4 40.5 50.5 51.7 42.9 54.4 46.4 54.2 48.6 53.1 50.1 35.1 40.0 34.6 40.1 39.9 49.7 48.3
max 55.9 50.8 46.2 48.6 52.5 41.6 42.2 54.2 55.8 46.0 56.8 50.3 58.2 52.6 56.0 54.0 37.8 42.6 37.0 43.5 43.9 51.4 51.9
mean,max 56.1 51.0 46.5 48.6 52.8 42.6 42.8 54.0 56.0 47.1 56.1 50.8 58.4 52.8 56.9 54.2 38.5 43.7 38.4 43.2 44.0 51.5 52.0
Table 5: Overall accuracy scores by LLM probe averaged over five runs training on the entire DISRPT (pooling strategy ablation). We ablate different attention-head-wise pooling strategies, namely mean pooling, maximum pooling, and the concatenation of both on the attention score matrices $\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}$. We find that concatenating both generally yields the best results, though usually only by a margin of less than $1\%$. To keep the size of the representations shorter, we thus opt to only keep the maximum pooling.